# A perturbative construction of primitive forms from log Landau-Ginzburg mirrors of toric manifolds

### v1

@@ Appendix A Dolbeault resolution
Take $\Delta$ to be the defining polytope of the projective toric manifold $Y$. Then each $p$-dimensional polytope $\sigma$ of $\Delta$ corresponds to an $(n-p)$-dimensional component of $X$, which is denoted by $\sigma^{\circ}$. Moreover, the inverse inclusion relation holds, i.e., $\sigma\subset\tau$ if and only if $\tau^{\circ}\subset\sigma^{\circ}$. For any $\sigma\in\Delta$, define
@@
The injection part is obvious. To prove the exactness, we should show that every $\delta^{0}$-closed $(g_{i})\in\mathcal{C}^{0}(\Delta,\mathscr{O})$ with $g_{i}\in\mathscr{O}_{X_{i}}$ is the restriction of a global function on $X$. By definition, $\delta^{0}(g_{i})=0$ implies $g_{i}|_{X_{i}\cap X_{j}}=g_{j}|_{X_{i}\cap X_{j}}$ if the intersection is $(n-1)$-dimensional. In fact, the same identity holds even if the intersection has codimension greater than $1$. This can be seen as follows. Since $X_{i}\cap X_{j}$ is a subset of both $X_{i}$ and $X_{j}$, both of the vertices $v_{i}$ and $v_{j}$ corresponding to $X_{i}$ and $X_{j}$ respectively are in the polytope $\sigma$ corresponding to $X_{i}\cap X_{j}$. We can find a sequence of vertices $v_{i},v_{s},v_{t}\cdots,v_{u},v_{j}$ of $\sigma$ such that each pair of adjacent vertices is connected by an edge in $\sigma$. Now by the definition of $\delta^{0}$-closedness, we have
@@
Let $\rho\in\Delta_{n-1}$ be an $(n-1)$-dimensional polytope and consider the complex $(\mathcal{C}(\rho,\tilde{\mathscr{O}}(\rho^{\circ})),\delta)$, where $\tilde{\mathscr{O}}(\rho^{\circ})$ denotes the space of functions on $\rho^{\circ}$ that restricts to 00 at $o$. Now closeness of $(g^{(1)}_{\sigma})$ implies the closeness of the restriction of $(g^{(1)}_{\sigma})_{\sigma\subset\rho}$ to a $p$-chain in $(\mathcal{C}(\rho,\tilde{\mathscr{O}}(\rho^{\circ})),\delta)$ and the latter can be viewed as a family of $p$-cycles in $(\mathcal{C}(\rho,\mathbb{C}),\delta)$ that are parametrized by points of $\rho^{\circ}$. By Lemma A.6, we can find a $\rho^{\circ}$-parametrized $(p-1)$-chain $(f^{\rho}_{\tau})_{\tau\in\rho_{p-1}}$ in $(\mathcal{C}(\rho,\mathbb{C}),\delta)$ such that
@@
Moreover, $(f^{\rho}_{\tau})_{\tau\in\rho_{p-1}}$ is a chain in $(\mathcal{C}(\rho,\tilde{\mathscr{O}}(\rho^{\circ})),\delta)$. If we define $f^{\rho}_{\tau}=0$ for $\tau\in\Delta_{p-1}\setminus\rho_{p-1}$, we get now a $(p-1)$-chain $(f^{\rho}_{\tau})_{\tau\in\Delta_{p-1}}$ in $(\mathcal{C}(\Delta,\tilde{\mathscr{O}}(\rho^{\circ})),\delta)$ and the projection $\tau^{\circ}\to\rho^{\circ}$ allows us to view it as a chain in $(\mathcal{C}(\Delta,\mathscr{O}),\delta)$. Define
@@
We claim that the restriction of $(g^{(2)}_{\sigma})$ to any $\rho^{\circ},\rho\in\Delta_{n-1}$ is zero. By construction, $((g^{(1)}_{\sigma})-\delta^{p-1}(f^{\rho}_{\tau}))$ restricts to zero on $\rho^{\circ}$. Given $\rho^{\prime}\in\Delta_{n-1},\rho^{\prime}\neq\rho$, if there exists $\tau\in\rho^{\prime}_{p-1}\cap\rho_{p-1}$, $f^{\rho^{\prime}}_{\tau}$ is the pull-back to $\tau^{\circ}$ of a function in $\tilde{\mathscr{O}}(\rho^{\prime\circ})$. As $\rho^{\prime}\neq\rho$, $\rho^{\circ}\cap\rho^{\prime\circ}$ is zero dimensional, hence by construction $f^{\rho^{\prime}}_{\tau}$ restrict to 00 on $\rho^{\circ}$. If no $(p-1)$-dimensional polytope is in $\rho^{\prime}_{p-1}\cap\rho_{p-1}$, then by definition, $f^{\rho^{\prime}}_{\tau}=0,\forall\tau\subset\rho$.
Inductively, we can define $(g^{(3)}_{\sigma}),(g^{(4)}_{\sigma}),\cdots,(g^{(n-p)}_{\sigma})\in(\mathcal{C}(\Delta,\mathscr{O}),\delta)$ such that: a) each $(g^{(i)}_{\sigma})$ differs from $(g^{(i-1)}_{\sigma})$ by a $\delta^{p-1}$-boundary; b) any function $g^{(i)}_{\sigma}$ restrict to 00 on $\rho^{\circ}$, where $\rho$ is any polytope that contains $\sigma$ and is of dimension not smaller than $n-i+1$.
Now $(g^{(n-p)}_{\sigma})_{\sigma\in\Delta_{p}}$ differs from $(g_{\sigma})_{\sigma\in\Delta_{p}}$ by a $\delta^{p-1}$-boundary and the lemma is proved if $(g^{(n-p)}_{\sigma})_{\sigma\in\Delta_{p}}$ can also be written as a $\delta^{p-1}$-boundary. Given any $g^{(n-p)}_{\sigma}$ on $\sigma^{\circ}$, let $\tau(\sigma)\subset\sigma$ be a $(p-1)$-dimensional polytope. The projection $p:\tau(\sigma)^{\circ}\to\sigma^{\circ}$ lifts $g^{(n-p)}_{\sigma}$ to a function on $\tau(\sigma)^{\circ}$, which is denoted by $g^{(n-p)}_{\tau(\sigma)}$. Since $g^{(n-p)}_{\sigma}$ restrict to 00 on $\rho^{\circ}$ for any $\rho$ properly contains $\sigma$, $g^{(n-p)}_{\tau(\sigma)}$ will restrict to 00 on any $p$ dimensional polytope $\sigma^{\prime}\in\Delta_{p}\setminus\{\sigma\}$ that contains $\tau(\sigma)$. Define a $(p-1)$-chain $(g^{(n-p)}_{\sigma;\tau})_{\tau\in\Delta_{p-1}}$ by letting $g^{n-p}_{\sigma;\tau(\sigma)}=g^{n-p}_{\tau(\sigma)}$ and $g^{(n-p)}_{\sigma;\tau^{\prime}}=0$ for any $\tau^{\prime}\in\Delta_{p-1}\setminus\{\tau(\sigma)\}$, then the $\sigma$-component of $\delta^{p-1}(g^{(n-p)}_{\sigma;\tau})$ is ±$\pm g^{(n-p)}_{\sigma}$, in which the sign depends on the orientations of $\sigma$ and $\tau$. We put the sign before $(g^{(n-p)}_{\sigma;\tau})$, which by abuse of notation we still denote by $(g^{(n-p)}_{\sigma;\tau})$, so that the $\sigma$-component of $\delta^{p-1}(g^{(n-p)}_{\sigma;\tau})_{\tau\in\Delta_{p-1}}$ is exactly $g^{(n-p)}_{\sigma}$. Taking summation over all $\sigma\in\Delta_{p-1}$, we finally have
@@ Appendix B Regularized integral
be the sheaf of meromorphic forms which are holomorphic on $M-D$ but possibly with arbitrary orders of poles along $D$. Denote
@@
By definition, $\omega\in\mathcal{A}^{k}(M,\star D)$ if and only if $\omega$ is smooth on $M-D$ and locally around any $p\in D$, $\omega$ is of the form
@@
Any $\omega\in\mathcal{A}^{p,\bullet}(M,\star D),p\geq 1$ can be written as
$\omega=\alpha+\partial\beta,\text{ where }\alpha\in\mathcal{A}^{p,\bullet}(M,\log D),\beta\in\mathcal{A}^{p-1,\bullet}(M,\star D).$
@@
[…]
@@
[…]
@@
[…]

### v3

@@ Appendix A Dolbeault resolution
Take $\Delta$ to be the defining polytope of the projective toric manifold $Y$. Then each $p$-dimensional polytope $\sigma$ of $\Delta$ corresponds to an (np)(n-p)-dimensional component of $X$, which is denoted by $\sigma^{\circ}$. Moreover, the inverse inclusion relation holds, i.e., $\sigma\subset\tau$ if and only if $\tau^{\circ}\subset\sigma^{\circ}$. For any $\sigma\in\Delta$, define
@@
The injection part is obvious. To prove the exactness, we should show that every $\delta^{0}$-closed $(g_{i})\in\mathcal{C}^{0}(\Delta,\mathscr{O})$ with $g_{i}\in\mathscr{O}_{X_{i}}$ is the restriction of a global function on $X$. By definition, $\delta^{0}(g_{i})=0$ implies $g_{i}|_{X_{i}\cap X_{j}}=g_{j}|_{X_{i}\cap X_{j}}$ if the intersection is (n1)(n-1)-dimensional. In fact, the same identity holds even if the intersection has codimension greater than $1$. This can be seen as follows. Since $X_{i}\cap X_{j}$ is a subset of both $X_{i}$ and $X_{j}$, both of the vertices $v_{i}$ and $v_{j}$ corresponding to $X_{i}$ and $X_{j}$ respectively are in the polytope $\sigma$ corresponding to $X_{i}\cap X_{j}$. We can find a sequence of vertices $v_{i},v_{s},v_{t}\cdots,v_{u},v_{j}$ of $\sigma$ such that each pair of adjacent vertices is connected by an edge in $\sigma$. Now by the definition of $\delta^{0}$-closedness, we have
@@
Let $\rho\in\Delta_{n-1}$ be an (n1)(n-1)-dimensional polytope and consider the complex $(\mathcal{C}(\rho,\tilde{\mathscr{O}}(\rho^{\circ})),\delta)$, where $\tilde{\mathscr{O}}(\rho^{\circ})$ denotes the space of functions on $\rho^{\circ}$ that restricts to 0 at $o$. Now closeness of $(g^{(1)}_{\sigma})$ implies the closeness of the restriction of $(g^{(1)}_{\sigma})_{\sigma\subset\rho}$ to a $p$-chain in $(\mathcal{C}(\rho,\tilde{\mathscr{O}}(\rho^{\circ})),\delta)$ and the latter can be viewed as a family of $p$-cycles in $(\mathcal{C}(\rho,\mathbb{C}),\delta)$ that are parametrized by points of $\rho^{\circ}$. By Lemma A.6, we can find a $\rho^{\circ}$-parametrized (p1)(p-1)-chain $(f^{\rho}_{\tau})_{\tau\in\rho_{p-1}}$ in $(\mathcal{C}(\rho,\mathbb{C}),\delta)$ such that
@@
Moreover, $(f^{\rho}_{\tau})_{\tau\in\rho_{p-1}}$ is a chain in $(\mathcal{C}(\rho,\tilde{\mathscr{O}}(\rho^{\circ})),\delta)$. If we define $f^{\rho}_{\tau}=0$ for $\tau\in\Delta_{p-1}\setminus\rho_{p-1}$, we get now a (p1)(p-1)-chain $(f^{\rho}_{\tau})_{\tau\in\Delta_{p-1}}$ in $(\mathcal{C}(\Delta,\tilde{\mathscr{O}}(\rho^{\circ})),\delta)$ and the projection $\tau^{\circ}\to\rho^{\circ}$ allows us to view it as a chain in $(\mathcal{C}(\Delta,\mathscr{O}),\delta)$. Define
@@
We claim that the restriction of $(g^{(2)}_{\sigma})$ to any $\rho^{\circ},\rho\in\Delta_{n-1}$ is zero. By construction, $((g^{(1)}_{\sigma})-\delta^{p-1}(f^{\rho}_{\tau}))$ restricts to zero on $\rho^{\circ}$. Given $\rho^{\prime}\in\Delta_{n-1},\rho^{\prime}\neq\rho$, if there exists $\tau\in\rho^{\prime}_{p-1}\cap\rho_{p-1}$, $f^{\rho^{\prime}}_{\tau}$ is the pull-back to $\tau^{\circ}$ of a function in $\tilde{\mathscr{O}}(\rho^{\prime\circ})$. As $\rho^{\prime}\neq\rho$, $\rho^{\circ}\cap\rho^{\prime\circ}$ is zero dimensional, hence by construction $f^{\rho^{\prime}}_{\tau}$ restrict to 0 on $\rho^{\circ}$. If no (p1)(p-1)-dimensional polytope is in $\rho^{\prime}_{p-1}\cap\rho_{p-1}$, then by definition, $f^{\rho^{\prime}}_{\tau}=0,\forall\tau\subset\rho$.
Inductively, we can define $(g^{(3)}_{\sigma}),(g^{(4)}_{\sigma}),\cdots,(g^{(n-p)}_{\sigma})\in(\mathcal{C}(\Delta,\mathscr{O}),\delta)$ such that: a) each $(g^{(i)}_{\sigma})$ differs from $(g^{(i-1)}_{\sigma})$ by a $\delta^{p-1}$-boundary; b) any function $g^{(i)}_{\sigma}$ restrict to 0 on $\rho^{\circ}$, where $\rho$ is any polytope that contains $\sigma$ and is of dimension not smaller than ni+1n-i+1.
Now $(g^{(n-p)}_{\sigma})_{\sigma\in\Delta_{p}}$ differs from $(g_{\sigma})_{\sigma\in\Delta_{p}}$ by a $\delta^{p-1}$-boundary and the lemma is proved if $(g^{(n-p)}_{\sigma})_{\sigma\in\Delta_{p}}$ can also be written as a $\delta^{p-1}$-boundary. Given any $g^{(n-p)}_{\sigma}$ on $\sigma^{\circ}$, let $\tau(\sigma)\subset\sigma$ be a (p1)(p-1)-dimensional polytope. The projection $p:\tau(\sigma)^{\circ}\to\sigma^{\circ}$ lifts $g^{(n-p)}_{\sigma}$ to a function on $\tau(\sigma)^{\circ}$, which is denoted by $g^{(n-p)}_{\tau(\sigma)}$. Since $g^{(n-p)}_{\sigma}$ restrict to 0 on $\rho^{\circ}$ for any $\rho$ properly contains $\sigma$, $g^{(n-p)}_{\tau(\sigma)}$ will restrict to 0 on any $p$ dimensional polytope $\sigma^{\prime}\in\Delta_{p}\setminus\{\sigma\}$ that contains $\tau(\sigma)$. Define a (p1)(p-1)-chain $(g^{(n-p)}_{\sigma;\tau})_{\tau\in\Delta_{p-1}}$ by letting $g^{n-p}_{\sigma;\tau(\sigma)}=g^{n-p}_{\tau(\sigma)}$ and $g^{(n-p)}_{\sigma;\tau^{\prime}}=0$ for any $\tau^{\prime}\in\Delta_{p-1}\setminus\{\tau(\sigma)\}$, then the $\sigma$-component of $\delta^{p-1}(g^{(n-p)}_{\sigma;\tau})$ is $\pm g^{(n-p)}_{\sigma}$, in which the sign depends on the orientations of $\sigma$ and $\tau$. We put the sign before $(g^{(n-p)}_{\sigma;\tau})$, which by abuse of notation we still denote by $(g^{(n-p)}_{\sigma;\tau})$, so that the $\sigma$-component of $\delta^{p-1}(g^{(n-p)}_{\sigma;\tau})_{\tau\in\Delta_{p-1}}$ is exactly $g^{(n-p)}_{\sigma}$. Taking summation over all $\sigma\in\Delta_{p-1}$, we finally have
@@ Appendix B Regularized integral
be the sheaf of meromorphic forms which are holomorphic on MDM-D but possibly with arbitrary orders of poles along $D$. Denote
@@
By definition, $\omega\in\mathcal{A}^{k}(M,\star D)$ if and only if $\omega$ is smooth on MDM-D and locally around any $p\in D$, $\omega$ is of the form
@@
Any $\omega\in\mathcal{A}^{n,\bullet}(M,\star D)$ can be written as
$\omega=\alpha+\partial\beta,\text{ where }\alpha\in\mathcal{A}^{p,\bullet}(M,\log D),\beta\in\mathcal{A}^{n-1,\bullet}(M,\star D).$
@@
$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{M}\omega:=\begin{cases}0&\text{if}\quad\omega\in\mathcal{A}^{<2n}(M,\star D)\,,\\ \int_{M}\alpha+\int_{\partial M}\beta&\text{if}\quad\omega=\alpha+\partial\beta\in\mathcal{A}^{2n}(M,\star D)\,.\end{cases}$
@@
$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{M}d\omega=-2\pi i\sum_{j=1}^{l}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{D_{j}}\operatorname{Res}_{D_{j}}(\omega)+\int_{\partial M}\omega\,.$
@@
$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{M}d\omega=-2\pi i\sum_{j=1}^{l}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{D_{j}}\operatorname{Res}_{D_{j}}(\omega)\,.$